# On the heat kernel of non-minimal second-order operators

### v1

@@ Abstract
We analyze the spectra of general non-minimal second-order operators. To do this, we derive the local part of the trace of the second Seeley-DeWitt heat kernel coefficient for such operators in a completely model-independent way. Afterwards, we provide three examples to show how our result can be applied in practical scenarios. In particular, we emphasize this discussion when dealing with a toy-model of dynamical torsion, which is viewed as a simple instance of higher-spin fields. All our results are compatible with the literature, and we provide a Mathematica notebook with the model-independent results that are written in the paper.
@@ I Introduction
The heat kernel method provides an elegant computational tool for deriving perturbative quantum properties in the presence of background fields in a manifestly covariant fashion. For this reason, it has been widely used in the context of gravitational theories (see, e.g.,Barvinsky:1985an ; Vassilevich:2003xt ; Percacci:2017fkn ). Among the many phenomenological applications we cite the computation of beta-functions Buchbinder:2021wzv ; Daas:2020dyo , working out the quantum anomalies Vassilevich:2005vk ; Barvinsky:2023exr , deriving the quantum corrections to the black hole entropy Fursaev:1994ea and analyzing gravitational catalysis Gies:2021upb . Furthermore, even though this formalism is mostly applied in the $1$-loop approximation, it can also be extended to take into account multi-loop diagrams Jack:1982hf ; Carneiro:2024slt ; Fuentes-Martin:2024agf ; Carneiro:2025xny .
Most of the applications of the heat kernel method have focused on minimal second-order operators, i.e., those whose principal part is given by a Laplace-type operator. This is because in this case one can prove that there exists an asymptotic expansion of the heat kernel for small distances Gilkey:1975iq , which in turn suggests the introduction of the so-called Seeley-DeWitt bi-tensors. Such an expansion was found in DeWitt:1964mxt by comparison with the known flat-space result of a diffusion process. Then, the coincidence limits of such coefficients can be found in a recursive way DeWitt:1964mxt . Of particular interest are the first three coefficients $\hat{a}_{0}$, $\hat{a}_{1}$ and $\hat{a}_{2}$, since these dictate the form of the beta-functions of marginal and relevant operators in $d=4$ in the $1$-loop approximation. In practice, the knowledge of the trace of $\hat{a}_{2}$ suffices to yield the beta-functions of dimensionless coupling constant; for this reason the focus of the present paper will be on the derivation of such a trace for a wide class of differential operators.
@@
On the other hand, if one wants to check the gauge invariance of a given calculation or considers more complicated fields such as torsion, non-metricity or higher-spins, he will have to deal with non-minimal operators. To derive the heat kernel coefficients of these operators a completely different approach is needed, since there exists no asymptotic expansion when the principal part is not of Laplace type. However, one can rely on a different perturbative expansion as in Barvinsky:1985an , or he can adopt the so-called non-local heat kernel Groh:2011dw . Recently, such a problem has been re-analyzed in some special examples Farolfi:2025knq ; Melichev:2025hcg ; Sangy:2025jxk . Of particular interest for the present discussion is the approach of Melichev:2025hcg , in which a slight modification of the trick introduced in Barvinsky:1985an was employed. Indeed, the approach of Melichev:2025hcg significantly improves the computational time needed for the derivation of the trace of the $\hat{a}_{2}$ coefficient. A radically different approach towards non-minimal operator is given by employing spin-parity decompositions, see, e.g. Groh:2011dw . Here the idea is to trade non-minimality for higher order but minimal operators. Such an approach was applied to a dynamical torsion theory in Martini:2023apm , but it will not be pursued here.
@@
Thus, the aim of this work is to derive local part of the trace of the second Seeley-DeWitt coefficient for general non-minimal operators in a model-independent way, and to provide some examples to show how the established recipe can be applied in practice. To this end the trick of Barvinsky:1985an is applied in the new formulation of Melichev:2025hcg , and vector bundles are employed to stand for an unspecified target space parameterizing the quantum fields of the theory. On a practical viewpoint the xAct packages of Mathematica Martin-Garcia:xAct ; Martin-Garcia:xTensor ; Martin-Garcia:2008ysv ; Nutma:2013zea have been heavily exploited, out of which FieldsX has played a crucial role Frob:2020gdh . Furthermore, the complete form of the model-independent result derived in this paper is also attached in a Mathematica notebook to provide a handy use of the present findings.
@@
The structure of the paper is as follows. First of all, in Sect. II the heat kernel is introduced, and it is related to the trace of the logarithm of a second-order Laplace operator. Next, in Section III non-minimal second order operators are considered in a formal way, singling out the trace-log of the principal part in Subsections III.1 and III.2 and treating perturbatively the rest in III.3. To turn formal expressions into concrete ones one has to evaluate functional traces, thus in Sect. IV a method to perform such calculations is presented. In Sect. V the final result for the trace-log of a general non-minimal second-order operator is written down. To this end, some computer-based computational tools and the associated conventions for representing field-space indices are briefly discussed in Subsect. V.1. Such conventions are employed in Subsect.s V.2 and V.3 to present the resulting modifications to the trace of the second Seeley-DeWitt coefficient $\hat{a}_{2}$. Section VI provides three examples of applications of the general findings, dealing with a vector field VI.1, a Kalb-Ramond field VI.2 and a toy-model of the torsion VI.3. The results are summarized and some possible future developments are discussed in Sect. VII. Finally, the Appendices contain useful commutator identities A, they provide all coincidence limits up to dimension-four terms B, and they collect the tensor structures that concur to ${\rm tr}\,\hat{a}_{2}$ that were omitted in Sect. V to avoid overburdening the flow of the main text C. Finally, the mostly-plus Minkowski signature will be used throughout the paper.
@@ II The Schwinger-DeWitt technique
$i\partial_{s}\langle x,s|y,0\rangle+\langle x|\hat{F}(\nabla)|0\rangle=0\,,\qquad\langle x,0|y,0\rangle=\delta(x,y)\,.$ (8)
For notational convenience, we shall denote the “transition amplitute” by
@@
which has to be thought of as a modification to the known flat-space solution in absence of background fields Schwinger:1951nm . Here and in the following $2\omega=d$, where $d$ is the spacetime dimension. Moreover, $\sigma(x,y)$ is the Synge world function, which is equal to half the square of the geodesic distance between $x$ and $y$. Thus, it thus generalizes the flat-space modulus of the difference of the $x$ and $y$ coordinates, and it obeys
@@
that describes the divergence or convergence of geodesics emanating from the point $x$, and it becomes singular whenever we encounter a caustic surface. Such a bi-scalar density can be re-written to single out the densities as
@@ III Non-minimal operators
III.1 Trace-log of the principal part: the simple case
@@
$\hat{D}(n;\zeta)\,\hat{K}(n;\zeta)=n^{2}\hat{\mathds{1}}\,$. (30)
Then, we promoting the vectors $n_{\mu}$ to covariant derivatives defines $\hat{K}(\hat{\nabla};\zeta)$ by fixing the arbitrary ordering of covariant derivatives. Then, due to the previous equation, we automatically find that
@@
where $\hat{M}$ is a differential operator of order two. Next, we compare Eq.s (28) and (31) to find a perturbative expression of $\hat{G}_{0}(\hat{\nabla};\zeta)$, which reads
@@
where we use ellipsis to stand for higher-dimensional operators that we are not interested in. To derive this equation we have commuted the inverse box operators to the right using Eq.s (98) and (100). Finally, we can insert Eq. (32) into Eq. (29), obtaining the desired relation for the trace-log of the principal part
@@
III.2 Trace-log of the principal part: the general case
@@
$\frac{d\hat{D}(\hat{\nabla};\zeta)}{d\zeta}\neq\hat{D}(\hat{\nabla};\zeta=1)\,.$ (34)
@@
where we have used the definition of the principal part. The expansion of the inverse $\hat{G}_{0}$ of $\hat{D}$ Eq. (32) can still be used. Thus, we can re-write the previous equation as
@@ III.3 Trace-log of the rest
The first term on the right-hand-side is given by Eq. (33) (or by Eq. (36) in the most general case), therefore we now focus on the second one. By using the known Taylor expansion of the logarithm and commuting the $\hat{\square}^{-2}$ operators to the right we find
@@
When the starting operator involves a derivative interaction $\hat{V}$, the method of Barvinsky:1985an requires the evaluation of expressions with up to $16$ covariant derivatives acting on $\hat{\square}^{-10}$. On the other hand, in the present case we must handle expressions with up to $12$ derivatives acting on $\hat{\square}^{-8}$. As we will show in the next section working out the precise form of these functional traces is not particularly involved. Instead, the computational cost increases with the number of derivatives, regardless the overall mass dimension of a given functional trace. This is due to the appearance of a number of symmetrized metric tensors which is at most equal to half the number of covariant derivatives. In general, handling these kind of expressions embodies a very expensive part of any such computation. Therefore, the present method significantly improves the computational efficiency of the calculation.
@@ IV Functional traces
which represent derivative interactions in loop diagrams with $n$ internal lines. Here we are only interested in the logarithmically divergent parts of these diagrams, which will be combined with the explicit form of the differential operators entering Eq.s (33) and (38).
The first step in the evaluation of Eq. (46) is provided by employing a Laplace transform to re-write the $\hat{\square}^{-n}$ operator as
@@
By inspecting of Eq.s (33), (36), (40), (41), (42), (43) and (44) we can list the functional traces that we need to evaluate
$\hat{\nabla}_{\mu_{1}}\hat{\nabla}_{\mu_{2}}\frac{\hat{\mathds{1}}}{\,\hat{\square}}\delta(x,x^{\prime})|^{\rm div}_{x^{\prime}=x}\,,\qquad\qquad\hat{\nabla}_{\mu_{1}}\hat{\nabla}_{\mu_{2}}\frac{\hat{\mathds{1}}}{\,\,\hat{\square}^{2}}\delta(x,x^{\prime})|^{\rm div}_{x^{\prime}=x}\,;\,;$ (53)
@@
Dimensional analysis already suggests that the structure of these objects may be written in an universal way, as we will show below. Let us start by considering 2n42n-4 covariant derivatives acting on the n-th inverse power of the box operator. By employing Eq. (50) we find
@@
$\hat{\nabla}_{\mu_{1}}\dots\hat{\nabla}_{\mu_{2n-2}}$ $\frac{\hat{\mathds{1}}}{\,\,\hat{\square}^{n}}\delta(x,x^{\prime})|^{\rm div}_{x^{\prime}=x}=-\frac{i\,\log L^{2}}{16\,\pi^{2}}\,\frac{\sqrt{-g}}{2^{n-2}\,(n-1)!}\left[\frac{1}{12}R\,g^{(n-1)}_{\mu_{1}\dots\mu_{2n-2}}\right$. (57)
$\left.-\sum_{i<j}\left(\frac{1}{6}R_{ij}+\frac{1}{2}\hat{\mathcal{R}}_{ij}\hat{\mathds{1}}\right)g^{(n-2)}_{\mu_{1}\dots\hat{i}\dots\hat{j}\dots\mu_{2n-2}}+\frac{1}{3}\sum_{{\scriptscriptstyle i<j<k<l}}\left(R_{kilj}+R_{likj}\right)g^{(n-3)}_{\mu_{1}\dots\hat{i}\dots\hat{l}\dots\mu_{2n-2}}\hat{\mathds{1}}\right$]\,.
Notice that we have employed our generalized definition of $g^{(k)}=0$ for any $k<0$, see Eq. (55), and we are using the shorthand $\hat{\nabla}_{i}=\hat{\nabla}_{\mu_{i}}$. By taking a further covariant derivative we end up with dimension-three terms on the right-hand-side, whose precise structure is
$\hat{\nabla}_{\mu_{1}}\dots\hat{\nabla}_{\mu_{2n-1}}$ $\frac{\hat{\mathds{1}}}{\,\,\hat{\square}^{n}}\delta(x,x^{\prime})|^{\rm div}_{x^{\prime}=x}=-\frac{i\,\log L^{2}}{16\,\pi^{2}}\,\frac{\sqrt{-g}}{2^{n-2}\,(n-1)!}\left[\frac{1}{12}\sum_{i}\left(\hat{\nabla}^{\alpha}\hat{\mathcal{R}}_{\mu_{i}\alpha}+\nabla_{\mu_{i}}R\right)g^{(n-1)}_{\mu_{1}\dots\hat{i}\dots\mu_{2n-1}}\right$.
$\left.-\frac{1}{3}\sum_{\scriptscriptstyle i<j<k}\left(\hat{\nabla}_{i}\hat{\mathcal{R}}_{jk}+\hat{\nabla}_{j}\hat{\mathcal{R}}_{ik}+\tfrac{1}{4}\nabla_{i}R_{jk}+\tfrac{1}{4}\nabla_{j}R_{ik}+\tfrac{1}{4}\nabla_{k}R_{ij}\right)g^{(n-2)}_{\mu_{1}\dots\hat{i}\dots\hat{k}\dots\mu_{2n-1}}\hat{\mathds{1}}\right$.
[…]
$\left.\left.+\nabla_{m}R_{iljk}\right)g^{(n-3)}_{\mu_{1}\dots\hat{i}\dots\hat{m}\dots\mu_{2n-1}}\hat{\mathds{1}}\right$]\,. (58)

### v2

# Heat kernel of non-minimal second-order operators

@@ Abstract
We analyze the spectra of general non-minimal second-order operators. To do this, we derive the local part of the trace of the second Seeley-DeWitt heat kernel coefficient for such operators in a completely model-independent way. Afterwards, we provide three examples to show how our result can be applied in practical scenarios. In particular, we emphasize this discussion when dealing with a toy model of dynamical torsion, which is viewed as a simple instance of higher-spin fields. All our results are compatible with the literature, and we provide a Mathematica notebook with the model-independent results that are written in the paper.
@@ I Introduction
The heat kernel method provides an elegant computational tool for deriving perturbative quantum properties in the presence of background fields in a manifestly covariant fashion. For this reason, it has been widely used in the context of gravitational theories (see, e.g.,Barvinsky:1985an ; Vassilevich:2003xt ; Percacci:2017fkn ). Among the many phenomenological applications we cite the computation of beta-functions Buchbinder:2021wzv ; Daas:2020dyo , working out the quantum anomalies Vassilevich:2005vk ; Barvinsky:2023exr , deriving the quantum corrections to the black hole entropy Fursaev:1994ea and analyzing gravitational catalysis Gies:2021upb . Furthermore, even though this formalism is mostly applied in the $1$-loop approximation, it can also be extended to take into account multi-loop diagrams Jack:1982hf ; Ivanov:2020yrc ; Ivanov:2024lbs ; Carneiro:2024slt ; Fuentes-Martin:2024agf ; Carneiro:2025xny .
Most of the applications of the heat kernel method have focused on minimal second-order operators, i.e., those whose principal part is given by a Laplace-type operator. This is because in this case one can prove that there exists an asymptotic expansion of the heat kernel for small distances Gilkey:1975iq , which in turn suggests the introduction of the so-called Seeley-DeWitt bi-tensors. Such an expansion was found in DeWitt:1964mxt by comparison with the known flat-space result of a diffusion process. Then, the coincidence limits of such coefficients can be found in a recursive way DeWitt:1964mxt . Of particular interest are the first three coefficients $\hat{a}_{0}$, $\hat{a}_{1}$ and $\hat{a}_{2}$, since these dictate the form of the beta-functions of marginal and relevant operators in $d=4$ in the $1$-loop approximation. In practice, the knowledge of the trace of $\hat{a}_{2}$ suffices to yield the beta-functions of dimensionless coupling constant; for this reason the focus of this paper will be on the derivation of such a trace for a wide class of differential operators.
@@
On the other hand, if one wants to check the gauge invariance of a given calculation or considers more complicated fields such as torsion, non-metricity or higher-spins, he will have to deal with non-minimal operators. To derive the heat kernel coefficients of these operators a completely different approach is needed, since there exists no asymptotic expansion when the principal part is not of Laplace type. However, one can rely on a different perturbative expansion as in Barvinsky:1985an , or he can adopt the so-called non-local heat kernel Groh:2011dw . Recently, such a problem has been reanalyzed in some special examples Grasso:2023qye ; Farolfi:2025knq ; Melichev:2025hcg ; Sangy:2025jxk . Of particular interest for the present discussion is the approach of Melichev:2025hcg , in which a slight modification of the trick introduced in Barvinsky:1985an was employed. Indeed, the approach of Melichev:2025hcg significantly improves the computational time needed for the derivation of the trace of the $\hat{a}_{2}$ coefficient. A radically different approach towards non-minimal operator is given by employing spin-parity decompositions, see, e.g. Groh:2011dw . Here the idea is to trade non-minimality for higher order but minimal operators. Such an approach was applied to a dynamical torsion theory in Martini:2023apm , but it will not be pursued here. Finally, non-minimal operators arise also in Proca-like theories, where the principal part is degenerate. In such cases one usually singles out the non-minimal part and derives the heat kernel coefficients by deriving the algebra of the pseudo-differential operators that appear in the Hessian Endo:1984sz ; Endo:1994yj ; Barvinsky:2025jbw .
@@
Thus, the aim of this work is to derive local part of the trace of the second Seeley-DeWitt coefficient for general non-minimal operators in a model-independent way, and to provide some examples to show how the established recipe can be applied in practice. To this end the trick of Barvinsky:1985an is applied in the new formulation of Melichev:2025hcg , and vector bundles are employed to stand for an unspecified target space parametrizing the quantum fields of the theory. On a practical viewpoint the xAct packages of Mathematica Martin-Garcia:xAct ; Martin-Garcia:xTensor ; Martin-Garcia:2008ysv ; Nutma:2013zea have been heavily exploited, out of which FieldsX has played a crucial role Frob:2020gdh . Furthermore, the complete form of the model-independent result derived in this paper is also attached in a Mathematica notebook to provide a handy use of the present findings Sauro:2025nb .
@@
The structure of the paper is as follows. First of all, in Sect. II the heat kernel is introduced, and it is related to the trace of the logarithm of a second-order Laplace operator. Next, in Section III non-minimal second order operators are considered in a formal way, singling out the trace-log of the principal part in Subsections III.1 and III.2 and treating perturbatively the rest in III.3. To turn formal expressions into concrete ones one has to evaluate functional traces, thus in Sect. IV a method to perform such calculations is presented. In Sect. V the final result for the trace-log of a general non-minimal second-order operator is written down. To this end, some computer-based computational tools and the associated conventions for representing field-space indices are briefly discussed in Subsect. V.1. Such conventions are employed in Subsect.s V.2 and V.3 to present the resulting modifications to the trace of the second Seeley-DeWitt coefficient $\hat{a}_{2}$. Section VI provides three examples of applications of the general findings, dealing with a vector field VI.1, a Kalb-Ramond field VI.2 and a toy model of the torsion VI.3. The results are summarized and some possible future developments are discussed in Sect. VII. Finally, the Appendices contain useful commutator identities A, they provide all coincidence limits up to dimension-four terms B, and they collect the tensor structures that concur to ${\rm tr}\,\hat{a}_{2}$ that were omitted in Sect. V to avoid overburdening the flow of the main text C. Finally, the mostly-plus Minkowski signature will be used throughout the paper.
@@ II The Schwinger-DeWitt technique
$i\partial_{s}\langle x,s|y,0\rangle+\langle x,s|\hat{F}(\nabla)|y,0\rangle=0\,,\qquad\langle x,0|y,0\rangle=\delta(x,y)\,.$ (8)
For notational convenience, we shall denote the “transition amplitude” by
@@
which has to be thought of as a modification to the known flat-space solution in absence of background fields Schwinger:1951nm . Here and in the following $2\omega=d$, where $d$ is the spacetime dimension. Moreover, $\sigma(x,y)$ is the Synge world function, which is equal to half the square of the geodesic distance between $x$ and $y$. Thus, it generalizes the flat-space modulus of the difference of the $x$ and $y$ coordinates, and it obeys
@@
that describes the divergence or convergence of geodesics emanating from the point $x$, and it becomes singular whenever we encounter a caustic surface. Such a bi-scalar density can be rewritten to single out the densities as
@@ III Non-minimal operators
III.1 Trace log of the principal part: The simple case
@@
$\hat{D}(n;\zeta)\,\hat{K}(n;\zeta)=n^{4}\hat{\mathds{1}}\,$. (30)
Then, promoting the vectors $n_{\mu}$ to covariant derivatives defines $\hat{K}(\hat{\nabla};\zeta)$ by fixing the arbitrary ordering of covariant derivatives. Next, due to the previous equation, we automatically find that
@@
where $\hat{M}$ is a differential operator of order two. Then, we compare Eqs. (28) and (31) to find a perturbative expression of $\hat{G}_{0}(\hat{\nabla};\zeta)$, which reads
@@
where we use ellipsis to stand for higher-dimensional operators that we are not interested in. To derive this equation we have commuted the inverse box operators to the right using Eqs. (98) and (100). Finally, we can insert Eq. (32) into Eq. (29), obtaining the desired relation for the trace-log of the principal part
@@
III.2 Trace log of the principal part: The general case
@@
$\frac{d\hat{D}(\hat{\nabla};\zeta)}{d\zeta}\neq\hat{D}(\hat{\nabla};\zeta=1)-\hat{\square}\,$. (34)
@@
where we have used the definition of the principal part. The expansion of the inverse $\hat{G}_{0}$ of $\hat{D}$ Eq. (32) can still be used. Thus, we can rewrite the previous equation as
@@ III.3 Trace-log of the rest
The first term on the right-hand side is given by Eq. (33) (or by Eq. (36) in the most general case), therefore we now focus on the second one. By using the known Taylor expansion of the logarithm and commuting the $\hat{\square}^{-2}$ operators to the right we find
@@
When the starting operator involves a derivative interaction $\hat{V}$, the method of Barvinsky:1985an requires the evaluation of expressions with up to $16$ covariant derivatives acting on $\hat{\square}^{-10}$. On the other hand, in the present case we must handle expressions with up to $12$ derivatives acting on $\hat{\square}^{-8}$. As we will show in the next section working out the precise form of these functional traces is not particularly involved. Instead, the computational cost increases with the number of derivatives, regardless the overall mass dimension of a given functional trace. This is due to the appearance of a number of symmetrized metric tensors which is at most equal to half the number of covariant derivatives. In general, handling these kinds of expressions embodies a very expensive part of any such computation. Therefore, the present method significantly improves the computational efficiency of the calculation.
@@ IV Functional traces
which represent derivative interactions in loop diagrams with $n$ internal lines. Here we are only interested in the logarithmically divergent parts of these diagrams, which will be combined with the explicit form of the differential operators entering Eqs. (33) and (38).
The first step in the evaluation of Eq. (46) is provided by employing a Laplace transform to rewrite the $\hat{\square}^{-n}$ operator as
@@
By inspecting of Eqs. (33), (36), (40), (41), (42), (43) and (44) we can list the functional traces that we need to evaluate
$\hat{\nabla}_{\mu_{1}}\hat{\nabla}_{\mu_{2}}\frac{\hat{\mathds{1}}}{\,\hat{\square}}\delta(x,x^{\prime})|^{\rm div}_{x^{\prime}=x}\,,\qquad\qquad\hat{\nabla}_{\mu_{1}}\hat{\nabla}_{\mu_{2}}\frac{\hat{\mathds{1}}}{\,\,\hat{\square}^{2}}\delta(x,x^{\prime})|^{\rm div}_{x^{\prime}=x}\,;$ (53)
@@
Dimensional analysis already suggests that the structure of these objects may be written in an universal way, as we will show below. Let us start by considering 2n42n-4 covariant derivatives acting on the $n$th inverse power of the box operator. By employing Eq. (50) we find
@@
$\hat{\nabla}_{\mu_{1}}\dots\hat{\nabla}_{\mu_{2n-2}}$ $\frac{\hat{\mathds{1}}}{\,\,\hat{\square}^{n}}\delta(x,x^{\prime})|^{\rm div}_{x^{\prime}=x}=-\frac{i\,\log L^{2}}{16\,\pi^{2}}\,\frac{\sqrt{-g}}{2^{n-2}\,(n-1)!}\left[\frac{1}{12}R\,g^{(n-1)}_{\mu_{1}\dots\mu_{2n-2}}\,\hat{\mathds{1}}\right$. (57)
$\left.-\sum_{i<j}\left(\frac{1}{6}R_{ij}\hat{\mathds{1}}+\frac{1}{2}\hat{\mathcal{R}}_{ij}\right)g^{(n-2)}_{\mu_{1}\dots\hat{i}\dots\hat{j}\dots\mu_{2n-2}}+\frac{1}{3}\sum_{{\scriptscriptstyle i<j<k<l}}\left(R_{kilj}+R_{likj}\right)g^{(n-3)}_{\mu_{1}\dots\hat{i}\dots\hat{l}\dots\mu_{2n-2}}\hat{\mathds{1}}\right$]\,.
Notice that we have employed our generalized definition of $g^{(k)}=0$ for any $k<0$, see Eq. (55), and we are using the shorthand $\hat{\nabla}_{i}=\hat{\nabla}_{\mu_{i}}$. By taking a further covariant derivative we end up with dimension-three terms on the right-hand side, whose precise structure is
$\hat{\nabla}_{\mu_{1}}\dots\hat{\nabla}_{\mu_{2n-1}}$ $\frac{\hat{\mathds{1}}}{\,\,\hat{\square}^{n}}\delta(x,x^{\prime})|^{\rm div}_{x^{\prime}=x}=-\frac{i\,\log L^{2}}{16\,\pi^{2}}\,\frac{\sqrt{-g}}{2^{n-2}\,(n-1)!}\left[\frac{1}{12}\sum_{i}\left(\hat{\nabla}^{\alpha}\hat{\mathcal{R}}_{\mu_{i}\alpha}+\frac{\hat{\mathds{1}}}{2}\nabla_{\mu_{i}}R\right)g^{(n-1)}_{\mu_{1}\dots\hat{i}\dots\mu_{2n-1}}\right$.
$\left.-\sum_{\scriptscriptstyle i<j<k}\left(\tfrac{2}{3}\hat{\nabla}_{(i}\hat{\mathcal{R}}_{j)k}+\tfrac{1}{4}\nabla_{(i}R_{jk)}\hat{\mathds{1}}\right)g^{(n-2)}_{\mu_{1}\dots\hat{i}\dots\hat{k}\dots\mu_{2n-1}}\right$. (58)
$\left.+\,\frac{1}{2}\sum_{{\scriptscriptstyle i<j<k<l<m}}\left(\nabla_{i}R_{(l|j|m)k}+\nabla_{j}R_{(l|i|m)k}+\nabla_{k}R_{(l|i|m)j}\right)g^{(n-3)}_{\mu_{1}\dots\hat{i}\dots\hat{m}\dots\mu_{2n-1}}\hat{\mathds{1}}\,\right$]\,.